OpenStudy (anonymous):

y = 2x^2 + ax + 14. a is a constant. it can also be written as y = 2(x - 3)^2 + b b is also a constant :P help me find the values of a and b. i'll give the next part once you find the first.

OpenStudy (anonymous):

1 minute

OpenStudy (anonymous):

ax-b+8=0

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

Given, y= 2x^2 + ax + 14, also y= 2(x-3)(x-3) + b

OpenStudy (anonymous):

So, 2x^2 + ax + 14 = 2(x-3)(x-3) + b

OpenStudy (anonymous):

Solve and u will get the above relation.

OpenStudy (anonymous):

what have you tried ?

OpenStudy (anonymous):

i'm trying to complete the square, for the first eq.

OpenStudy (anonymous):

Hm no need just expand the second and compare the coefficients, you will get \( a=-12 \) and \( b=-4 \)

OpenStudy (anonymous):

can u explain in detail? How u got the coefficient?

OpenStudy (anonymous):

\[ 2x^2 + ax + 14 = 2(x - 3)^2 + b \Rightarrow 2x^2 + ax + 14 = 18-12 x+2 x^2 +b \] Now compare the coefficients of both sides, to get the result.

OpenStudy (anonymous):

Is this extended maths?

OpenStudy (anonymous):

Plzz say the last step

OpenStudy (anonymous):

ohhhhhh, smart. I never thought of that :P nopeee, this is add maths.

OpenStudy (anonymous):

Now, ax + 14= b+18-12x

OpenStudy (anonymous):

\[ 2x^2 + ax + 14 = 2 x^2 -12 x+ 18 +b \] Now ?

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

Think adarsh it's not much difficult after this..

OpenStudy (anonymous):

and add maths is advance maths?

OpenStudy (anonymous):

see, i got like this: ax - b= -12x + 4

OpenStudy (anonymous):

yes so, that means a=12 and b=-4.

OpenStudy (anonymous):

Ohhhhhhhhhhhhhh, I understand. THanksssssssssss.

OpenStudy (anonymous):

Glad to help.

OpenStudy (anonymous):

Thanbks a lot, FOOLForMaths. And Thyanks Tanvi, for bringing such a challenging question.

OpenStudy (anonymous):

y = 2x^2 + ax + 14 y = 2(x - 3)^2 + b \[=> 2x^2 + ax + 14 = 2(x^2-6x+9)+b\] \[=>2x^2 + ax + 14 = 2x^2-12x+(18+b)\] equating on both sides, we get \[a = -12x; 14=18+b\] so, a = -6 and b = -4

OpenStudy (anonymous):

^^ a=-6 ?!